# ThinkNote: Enhancing Knowledge Integration and Utilization of Large Language Models via Constructivist Cognition Modeling

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated strong performance across a wide range of NLP tasks. However, they often exhibit suboptimal behaviors and inconsistencies when exposed to unfamiliar external information, underscoring their limitations in effectively leveraging such knowledge. Inspired by constructivist learning theory, we propose ThinkNote, a novel framework that enhances the external knowledge utilization of LLMs through a two-stage constructivist cognitive modeling process. Specifically, ThinkNote performs knowledge assimilation to align new information with the model’s parametric memory, forming a coherent internal representation. It then applies thought accommodation to adapt internal reasoning, thereby promoting more consistent and reliable outputs. Extensive experimental results demonstrate that ThinkNote achieves a 10% improvement over strong baseline methods on various question-answering benchmarks. Further analysis indicates that ThinkNote effectively integrates and utilizes external knowledge to help LLMs generate accurate responses and improves their self-consistency. All data and codes are available at https://github.com/OpenMatch/ThinkNote.
@@
Zhipeng Xu1, Zhenghao Liu1thanks: indicates corresponding author., Yukun Yan2, Shuo Wang2, Shi Yu2 Zheni Zeng2, Chaojun Xiao2, Zhiyuan Liu2, Ge Yu1 and Chenyan Xiong3 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
@@ 1 Introduction
Large Language Models (LLMs), e.g. GPT-4 OpenAI (2023) and LLaMA Touvron et al. (2023), have shown strong emergent abilities and achieved convincing performance in various NLP tasks Wei et al. (2022a); Zhao et al. (2023a). Despite these advancements, LLMs often face challenges in effectively leveraging external knowledge to generate accurate responses, especially when encountering information that is noisy or incomplete Liu et al. (2023); Xie et al. (2024); Asai et al. (2023). Such limitations lead to response inconsistencies and hallucinations, constraining their application in knowledge-intensive scenarios Ji et al. (2023); Xu et al. (2024a); Wang et al. (2023).
Recent research investigates strategies for enhancing the integration and utilization of external knowledge in LLMs. These efforts aim to develop specialized modules that refine, reflect on, or summarize external knowledge before integration Yu et al. (2023a); Xu et al. (2023) or filter the external knowledge based on its relevance to the current context Peng et al. (2023); Zhao et al. (2023b); Asai et al. (2023). While these methods show their effectiveness in helping LLMs acquire new knowledge Ji et al. (2023); Wei et al. (2024), they still treat LLMs as passive recipients of information Steffe and Gale (1995); Larochelle et al. (1998) and fail to address the underlying cognitive limitations that lead to insufficient knowledge utilization Yu et al. (2025). This limitation leads to incorrect responses and even causes degenerated responses on certain tasks Foulds et al. (2024); Shuster et al. (2021).
@@
Our experiments in various knowledge-intensive question-answering tasks demonstrate the effectiveness of ThinkNote. Compared to utilizing the external knowledge directly, ThinkNote achieves a 10% improvement regardless of the scale of the LLMs used as backbone models. Our further analysis shows that ThinkNote causes information accumulation Fan et al. (2025) on specific tokens, suggesting that ThinkNote can effectively utilize critical elements of external information in generating responses.
@@ 2 Related Work
To mitigate these limitations, some methods attempt to enhance knowledge integration by optimizing the interface between external information and the model’s input context Cuconasu et al. (2024); Gao et al. (2023b). These methods typically filter or reformat noisy inputs through knowledge summarization Xu et al. (2023), reflection Asai et al. (2023), or multi-step verification Yu et al. (2023a); Peng et al. (2023). While these approaches have advanced the ability of LLMs to integrate and utilize external information, they often push the burden of knowledge alignment to external modules Wang et al. (2024), where the LLMs remain conceptualized as passive processors, failing to reorganize or evaluate new information based on their current understanding Madaan et al. (2023).
More recent efforts explore internal modifications to how LLMs represent, revise, or adapt knowledge. These include work on memory editing Meng et al. (2022); Trivedi et al. (2022), belief tracking Wilie et al. (2024), and metacognitive strategies such as self-verification and response calibration Madaan et al. (2023); Liu et al. (2023). While promising, these approaches are often fragmented and tool-like: they provide mechanisms to intervene in model behavior without modeling the interaction between the model and external knowledge Collins et al. (2024). There is still limited consensus on what constitutes effective knowledge integration at the cognitive level, and most current methods operate under engineering-driven constraints rather than principles drawn from learning theory or epistemology Banino et al. (2023); Floridi and Nobre (2024); Van Rooij et al. (2024).
@@ 3 Methodology
Figure 1: The Overview of the ThinkNote Workflow.
@@ 4 Experimental Methodology
RAG Models. We implement three widely used RAG models as baselines, including vanilla RAG, Chain-of-Note Yu et al. (2023a), and Self-RAG Asai et al. (2023). The vanilla RAG model directly feeds retrieved passages as the context and asks the LLM to generate the answer Ram et al. (2023a). Chain-of-Note extends the CoT method to summarize the query-related knowledge from external knowledge for producing answers. Self-RAG Asai et al. (2023) trains LLMs to employ a self-reflection mechanism to filter out irrelevant evidence effectively.
@@ 5.1 Overall Performance
Compared to vanilla RAG models, ThinkNote delivers over a 10% improvement in performance, demonstrating its effectiveness in enabling LLMs to more accurately integrate and utilize external knowledge for question answering. While self-refinement methods focus on denoising external information to enhance generation, ThinkNote surpasses these approaches with over a 4% improvement, highlighting the importance of fostering a deeper understanding of knowledge within LLMs rather than merely extracting surface-level content related to queries. Furthermore, ThinkNote consistently outperforms all baseline models, regardless of the model scale or whether the backbone is black or white-box, indicating its strong generalization capabilities.
@@
[…]
@@ 5.3 Information Flow Analysis for ThinkNote in Knowledge Utilization
This section focuses on analyzing how knowledge assimilation and thought accommodation influence the reasoning process. Specifically, we randomly sample a case from the NQ dataset and follow Fan et al. (2025) to utilize the saliency score to examine interactions between tokens, which is also referred to as the information flow. Further details on the saliency score are provided in Appendix A.2.
@@
Table 3: Experimental Results on Evaluating the Knowledge Usage Ability of Different RAG Models.
@@ 5.4 Effectiveness of ThinkNote in Knowledge Integration and Utilization
In this subsection, we evaluate the capability of ThinkNote to accurately utilize external knowledge to generate responses.
We evaluate the sensitivity of both the vanilla LLM and ThinkNote to noisy external knowledge by feeding them the top-$K$ retrieved documents. Figures LABEL:fig:popqars and LABEL:fig:nqrs show the top-$K$ retrieval accuracy, which reflects the concentration of query-relevant information in documents. As $K$ increases from 1 to 5, retrieval accuracy consistently improves, indicating a higher likelihood of relevant content being included. However, this improvement saturates after $K=3$, showing that additional documents tend to introduce more noise than useful information. Figures LABEL:fig:popqa and LABEL:fig:nq compare the performance of ThinkNote and baseline methods. While the performance of Vanilla RAG remains stable or even degrades when $K\geq 3$, indicating its limited robustness to noisy inputs, ThinkNote continues to benefit from the additional documents. This trend highlights ThinkNote’s superior ability to identify and utilize query-relevant information even in the presence of increased noise.
@@ 5.5 Effectiveness of ThinkNote in Improving Self-Consistency of LLMs
In this subsection, we explore the self-consistency of ThinkNote and baseline models in answering questions. Specifically, we conduct 100 sampling iterations to calculate the ratio of correction. The accuracy reflects the model’s tendency to consistently generate correct or incorrect answers. We compare three methods: Direct QA, Vanilla RAG, and ThinkNote.
We first show the accuracy of ThinkNote and baseline models in Figure LABEL:fig:consitency_acc. ThinkNote demonstrates its effectiveness by achieving a median accuracy of around 47%, outperforming the Vanilla RAG (10%) and the Direct QA (31%). Notably, the much lower median accuracy of the Vanilla RAG compared to the Direct QA suggests that retrieved documents may negatively influence LLMs’ ability to answer certain questions Xie et al. (2024). Figure LABEL:fig:consitency_density further illustrates the density distribution of answering accuracy. Compared with the Vanilla RAG, our ThinkNote improves self-consistency by reducing the density of examples with accuracy between 30% and 80%. Different from Vanilla RAG, ThinkNote increases the number of queries answered correctly, highlighting its capacity to help LLMs effectively leverage external knowledge.
@@ A.2 Further Details on the Computation of Saliency Scores
where $A_{h,l}$ denotes the attention matrix of the $h$-th head in layer $l$, $L(x)$ is the loss function (cross-entropy), and $\odot$ represents the Hadamard product. This formulation combines attention weights with their gradient-based sensitivities, allowing us to capture both what the model focuses on and how crucial that focus is to the final prediction. We obtain $A_{h,l}$ from layer 21 in our experiments, using Meta-Llama-3-8B-Instruct as the foundation model.
A.3 Case Study

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated strong performance across a wide range of NLP tasks. However, they often exhibit suboptimal behaviors and inconsistencies when exposed to unfamiliar external information, underscoring their limitations in effectively leveraging such knowledge. Inspired by constructivist learning theory, we propose ThinkNote, a novel framework that enhances the external knowledge utilization of LLMs through a two-stage constructivist cognitive modeling process. Specifically, ThinkNote performs knowledge assimilation to align new information with the model’s parametric memory, forming a coherent internal representation. It then applies thought accommodation to adapt internal reasoning, thereby promoting more consistent and reliable outputs. Extensive experimental results demonstrate that ThinkNote achieves a 10% improvement over strong baseline methods on various question-answering benchmarks. Further analysis indicates that ThinkNote effectively integrates and utilizes external knowledge to help LLMs generate accurate responses and improve their self-consistency. All data and codes are available at https://github.com/OpenMatch/ThinkNote.
@@
Zhipeng Xu1, Zhenghao Liu1thanks: indicates corresponding author., Yukun Yan2, Shuo Wang2, Shi Yu2, Zheni Zeng2, Chaojun Xiao2, Zhiyuan Liu2, Ge Yu1 and Chenyan Xiong3 1School of Computer Science and Engineering, Northeastern University, Shenyang, China 2Department of Computer Science and Technology, Tsinghua University, Beijing, China 3Language Technologies Institute, Carnegie Mellon University, Pittsburgh, United States
@@ 1 Introduction
Large Language Models (LLMs), e.g. ChatGPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have shown strong emergent abilities and achieved convincing performance in various NLP tasks Wei et al. (2022a); Zhao et al. (2023a). Despite these advancements, LLMs often face challenges in effectively leveraging external knowledge to generate accurate responses, especially when encountering information that is noisy or incomplete Liu et al. (2024); Xie et al. (2024); Asai et al. (2023). Such limitations lead to response inconsistencies and hallucinations, constraining their application in knowledge-intensive scenarios Ji et al. (2023); Xu et al. (2024a); Wang et al. (2023).
Recent research investigates strategies for enhancing the integration and utilization of external knowledge in LLMs. These efforts aim to develop specialized modules that refine, reflect on, or summarize external knowledge before integration Yu et al. (2023a); Xu et al. (2023) or filter the external knowledge based on its relevance to the current context Peng et al. (2023); Zhao et al. (2023b); Asai et al. (2023). While these methods show their effectiveness in helping LLMs acquire new knowledge Ji et al. (2023); Wei et al. (2024), they still treat LLMs as passive recipients of information Steffe and Gale (1995); Larochelle et al. (1998) and fail to address the underlying cognitive limitations that lead to insufficient knowledge utilization Yu et al. (2025). This limitation leads to incorrect responses and can even cause degraded responses on certain tasks Foulds et al. (2024); Shuster et al. (2021).
@@
Our experiments on a range of knowledge-intensive question-answering tasks demonstrate the effectiveness of ThinkNote. Compared with directly incorporating external knowledge, ThinkNote consistently achieves around a 10% performance improvement across backbone LLMs of different scales. Moreover, ThinkNote exhibits stronger robustness when facing noisy, incomplete, or partially misleading external knowledge, suggesting that actively constructing and reorganizing knowledge is more reliable than passive conditioning on retrieved evidence. Further analysis reveals that ThinkNote leads to more concentrated information accumulation Fan et al. (2025) on a small set of salient tokens, indicating its ability to identify and leverage critical elements of external knowledge during response generation.
@@ 2 Related Work
To mitigate these limitations, some methods attempt to enhance knowledge integration by optimizing the interface between external information and the model’s input context Cuconasu et al. (2024); Gao et al. (2023b). These methods typically aim to filter or reformat noisy inputs through knowledge summarization Xu et al. (2023), reflection Asai et al. (2023), or multi-step verification Yu et al. (2023a); Peng et al. (2023). While these approaches have advanced the ability of LLMs to integrate and utilize external information, they often push the burden of knowledge alignment to external modules Wang et al. (2025), where the LLMs remain conceptualized as passive processors, failing to reorganize or evaluate new information based on their current understanding Madaan et al. (2023). This design choice limits their ability to support adaptive knowledge integration, especially in scenarios that require the model to actively construct and revise its understanding.
More recent efforts explore internal modifications to how LLMs represent, revise, or adapt knowledge. These include work on memory editing Meng et al. (2022); Trivedi et al. (2023), belief tracking Wilie et al. (2024), and metacognitive strategies such as self-verification and response calibration Madaan et al. (2023); Liu et al. (2024). While promising, these approaches are often fragmented and tool-like: they provide mechanisms to intervene in model behavior without modeling the interaction between the model and external knowledge Collins et al. (2024). There is still limited consensus on what constitutes effective knowledge integration at the cognitive level, and most current methods operate under engineering-driven constraints rather than principles drawn from learning theory or epistemology Floridi and Nobre (2024).
@@ 3 Methodology
Figure 1: The Overview of the ThinkNote Workflow. The framework follows a two-stage process of knowledge assimilation and thought accommodation to integrate external knowledge into model reasoning.
@@ 4 Experimental Methodology
RAG Models. We implement three widely used RAG models as baselines, including vanilla RAG, Chain-of-Note Yu et al. (2023a), and Self-RAG Asai et al. (2023). The vanilla RAG model directly feeds retrieved passages as the context and asks the LLM to generate the answer Ram et al. (2023). Chain-of-Note extends the CoT method to summarize the query-related knowledge from external knowledge for producing answers. Self-RAG Asai et al. (2023) trains LLMs to employ a self-reflection mechanism to filter out irrelevant evidence effectively.
@@ 5.1 Overall Performance
Compared to vanilla RAG models, ThinkNote delivers over a 10% improvement in performance, demonstrating its effectiveness in enabling LLMs to more accurately integrate and utilize external knowledge for question answering. While self-refinement methods focus on denoising external information to enhance generation, ThinkNote surpasses these approaches with over a 4% improvement, highlighting the importance of fostering a deeper understanding of knowledge within LLMs rather than merely extracting surface-level content related to queries. Furthermore, ThinkNote consistently outperforms all baseline models, regardless of the model scale or whether the backbone is black or white-box, indicating strong generalization and robustness across different model configurations.
@@
Table 2: Ablation Study of ThinkNote. We examine the contributions of knowledge assimilation and thought accommodation, and vary instruction designs to simulate different learning behaviors. Experiments are conducted with Meta-Llama-3-Instruct-8B and Meta-Llama-3-Instruct-70B.
@@ 5.3 Information Flow Analysis for ThinkNote in Knowledge Utilization
We further analyze how knowledge assimilation and thought accommodation influence the reasoning process in Figure 5(a). Specifically, we sample a case from the NQ dataset and follow Fan et al. (2025) in utilizing the saliency score to examine interactions between tokens, which is also referred to as information flow. Further details on the saliency score are provided in Appendix A.2.
@@
Table 3: Experimental Results on Evaluating the Knowledge Usage Ability of Different RAG Models. Results are reported under three scenarios, evaluating the capabilities of models in external knowledge utilization, robustness to missing evidence, and knowledge conflict resolution.
@@ 5.4 Effectiveness of ThinkNote in Knowledge Integration and Utilization
In this subsection, we evaluate the capability of ThinkNote to utilize external knowledge to generate responses accurately.
As shown in Figure 5(b), we evaluate the sensitivity of both the vanilla RAG and ThinkNote to noisy external knowledge by feeding them the top-$K$ retrieved documents. Figures LABEL:fig:popqars and LABEL:fig:nqrs show the top-$K$ retrieval accuracy, which reflects the concentration of query-relevant information in documents. As $K$ increases from 1 to 5, retrieval accuracy consistently improves, indicating a higher likelihood of relevant content being included. However, this improvement saturates after $K=3$, showing that additional documents tend to introduce more noise than useful information. Figures LABEL:fig:popqa and LABEL:fig:nq compare the performance of ThinkNote and baseline methods. While the performance of Vanilla RAG remains stable or even degrades when $K\geq 3$, indicating its limited robustness to noisy inputs, ThinkNote continues to benefit from the additional documents. This trend highlights ThinkNote’s superior ability to identify and utilize query-relevant information even in the presence of increased noise.
@@ 5.5 Effectiveness of ThinkNote in Improving Self-Consistency of LLMs
In this subsection, we explore the self-consistency of ThinkNote and baseline models in answering questions, as shown in Figure 7(a). Specifically, we conduct 100 sampling iterations to calculate the ratio of correction. The accuracy reflects the model’s tendency to generate correct or incorrect answers consistently. We compare three methods: Direct QA, Vanilla RAG, and ThinkNote. For Vanilla RAG and Direct QA, the answer generation module is used for sampling, while ThinkNote utilizes the thought accommodation to generate sampled responses.
@@
We first show the accuracy of ThinkNote and baseline models in Figure LABEL:fig:consitency_acc. ThinkNote demonstrates its effectiveness by achieving a median accuracy of around 47%, outperforming the Vanilla RAG (10%) and the Direct QA (31%). Notably, the substantially lower median accuracy of Vanilla RAG compared to Direct QA indicates that incorporating retrieved documents can hinder the LLM’s ability to produce correct answers, rather than improve it Xie et al. (2024).
Figure LABEL:fig:consitency_density further illustrates the density distribution of answering accuracy. Compared with the Vanilla RAG, our ThinkNote improves self-consistency by reducing the density of examples with accuracy between 30% and 80%. In contrast to vanilla RAG, ThinkNote substantially increases the number of queries answered correctly, highlighting its stronger capability to guide LLMs in effectively leveraging external knowledge to produce reliable answers.
@@ 6 Conclusion
Acknowledgments
This work is partly supported by the National Natural Science Foundation of China (No. 62576082 and No. 62461146205) and CCF-Zhipu Large Model Innovation Fund (No. 202403).
@@ A.2 Further Details on the Computation of Saliency Scores
where $A_{h,l}$ denotes the attention matrix of the $h$-th head in layer $l$, $L(x)$ is the loss function (cross-entropy), and $\odot$ represents the Hadamard product. This formulation combines attention weights with their gradient-based sensitivities, allowing us to capture both what the model focuses on and how crucial that focus is to the final prediction. We extract $A_{h,l}$ from the 21st layer, using Meta-Llama-3-8B-Instruct as the foundation model.
A.3 More Details of Baseline Models
[…]
A.4 Case Study